OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)?

OpenStudy (anonymous):

First we find the slope: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{8-2}{4-6}\) Subtract & then divide

OpenStudy (anonymous):

6/-2

OpenStudy (anonymous):

=2/6

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yes, what's 6 / -2?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

am i right @iGreen.

OpenStudy (anonymous):

Yes, so the slope is -3, now let's plug this in with another point in point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - 2 = -3(x - 6)\) Distribute -3 into the parenthesis: \(y - 2 = -3x + 18\) Now add 2 to both sides, what's 18 + 2?

OpenStudy (anonymous):

20